### v1

# Two questions on Kneser colorings

Abstract
In this paper, we investigate two questions on Kneser graphs $KG_{n,k}$. First, we prove that the union of $s$ non-trivial intersecting families in (${[n]\choose k}$ has size at most (${n\choose k}-{n-s\choose k}$ for all sufficiently large $n$ that satisfy $n>(2+\epsilon)k^{2}$ with $\epsilon>0$. We provide an example that shows that this result is essentially tight for the number of colors close to $\chi(KG_{n,k})=n-2k+2$. We also improve the result of Bulankina and Kupavskii on the choice chromatic number, showing that it is at least $\frac{1}{16}n\log n$ for all $k<\sqrt{n}$ and $n$ sufficiently large.
1 Introduction
For positive integers $n,k$, let $[n]:=\{1,\dots,n\}$ denote the standard $n$-element set and let (${[n]\choose k}:=\{F\subset[n]:|F|=k\}$ be the set of its $k$-element subsets. For a set $X$ we denote its power set by $2^{X}$. A family is a collection of sets $\mathcal{F}\subset 2^{X}$. We say that a family $\mathcal{F}$ is intersecting if for any $A,B\subset\mathcal{F}$, we have $A\cap B\neq\emptyset$. One of the classical results in extremal combinatorics is the Erdős–Ko–Rado theorem [4], which states that for $n\geq 2k$ the largest intersecting family in (${[n]\choose k}$ has size at most (${n-1\choose k-1}$. For $n>2k$ equality holds only if $\mathcal{F}$ is of the form $\mathcal{S}_{x}:=\{F\subset{[n]\choose k}:x\in F\}$, for some $x\in[n]$. The covering number $\tau(\mathcal{F})$ of $\mathcal{F}$ is the size of the smallest set $X$ such that $X\cap F\neq\emptyset$ for all $F\in\mathcal{F}$. Then $\mathcal{S}_{x}$ has covering number $1$. Such families are called stars or trivial, and $x$ is called their center. Hilton and Milner [8] gave a tight upper bound on the size of intersecting family in (${[n]\choose k}$ with covering number at least $2$ for $n>2k$:
||($|\mathcal{F}|\leq{n-1\choose k-1}-{n-k-1\choose k-1}+1.$ (1.1)
The extremal families have the property that all but one ‘exceptional’ set can be pierced by a single element.
The following generalization of the Erdős–Ko–Rado theorem was proposed by Erdős [3]: for given $n,k,s$, determine, what is the largest size of a family ($\mathcal{F}\subset{[n]\choose k}$ such that $\mathcal{F}$ is a union of at most $s$ intersecting families. It is a natural conjecture to make that for triples of parameters $(n,k,s)$ where $n$ is somewhat large in comparison to $k,s$, the extremal example is a union of $s$ stars. Erdős indeed conjectured this for all $n\geq 2k+t-1$. In this case, ||=($|\mathcal{F}|={n\choose k}-{n-s\choose k}$. Frankl and Fűredi [5] showed that the conjecture is true for $s=2$ and $n>ck$ with $c\approx 2.62$. However, it is not true in general. This was shown for $n=2k+t-1$ by Hilton and Füredi. Frankl and Füredi also gave an example which shows that the extreme is different even for $s=2$, when $n\leq 2k+c_{0}\sqrt{k}$ for $n\geq n_{0}$ and some positive constant $c_{0}$. Ellis and Lifshitz [2] recently proved that for constant $s$ and $n>2k+C(s)k^{2/3}$ the union of stars is optimal, where $C(s)$ is a function of $s$.
It is very natural to interpret this problem in terms of Kneser graphs. Consider a Kneser graph $KG_{n,k}$, whose vertex set is (${[n]\choose k}$, and edges connect pairs of disjoint sets. Kneser showed that $\chi(KG_{n,k})\leq n-2k+2$ and conjectured that this is tight. Lovász proved [10] that this indeed holds. By now, there is a series of results on the chromatic number of Kneser graphs and different subgraphs (e.g., see the recent paper of Kaiser and Stehlík [11], where they construct edge-critical subgraphs of Kneser graphs with the same chromatic number). However, all these proofs are topological and thus are difficult to adapt to other similar settings. The relation to the above question is not difficult to discern. Actually, intersecting families in (${[n]\choose k}$ are exactly the independent sets in $KG_{n,k},$ and thus are the collections of vertices that can be colored by the same color. In these terms, the question of Erdős may be reformulated as follows: what is the maximum number of vertices of $KG_{n,k}$ that can be properly colored in $s$ colors?
In this paper, we answer this question in the regime when the number of colors can be very large, but $n$ is somewhat large compared to $k$.
Theorem 1.
Let $\epsilon>0$ and let $n,k$ be integers such that $n>(2+\epsilon)k^{2}$ and $n$ is sufficiently large. If $\mathcal{F}$ is a union of $s$ non-trivial intersecting families, $s<n$, then ||<($|\mathcal{F}|<{n\choose k}-{n-s\choose k}$.
A simple argument using the Hilton–Milner theorem gives that the conclusion of the theorem is valid for, roughly speaking, $n\geq 2k^{3}$. We present it in the next section.
Why do we consider the case when all families are non-trivial? If $\mathcal{F}_{1},\dots,\mathcal{F}_{s}\subset 2^{[n]}$ are $s$ different stars, then clearly $|\cup_{i\in[s]}\mathcal{F}_{i}|={n\choose k}-{n-s\choose k}$. Now let $\mathcal{F}_{1}$ be a star with the center in $1$. Remove $1$ from the ground set, also removing all sets which contain $1$, and $\mathcal{F}_{1}$ in particular. The problem for $(n,k,s)$ is reduced to the problem for $(n-1,k,s-1)$. Assuming that we have $|\cup_{i=2}^{s}\mathcal{F}_{i}|\leq{n-1\choose k}-{n-1-(s-1)\choose k}={n-1\choose k}-{n-s\choose k}$, we get $|\cup_{i\in[s]}\mathcal{F}_{i}|\leq{n-1\choose k-1}+{n-1\choose k}-{n-s\choose k}={n\choose k}-{n-s\choose k}$. If we get rid of the assumption of all families being non-trivial then we get the following corollary.
Corollary 1.
Let $\epsilon>0$ and let $n,s,k$ be integers such that $n>(2+\epsilon)k^{2}+s$ and $n$ is sufficiently large. If $\mathcal{F}$ is a union of $s$ intersecting families then ||($|\mathcal{F}|\leq{n\choose k}-{n-s\choose k}$. Equality is possible only if $\mathcal{F}$ is a union of $s$ stars.
We also show that the bounds in Theorem 1 are not far from being tight.
Theorem 2.
Let $n=(k-1)(k+1-m)+m$, where $k,m$ are positive integers and $m\leq\frac{k}{4}$ and $k$ is sufficiently large. Then there exists a collection of $s:=n-2k+2-m$ non-trivial intersecting families in (${[n]\choose k}$ such that their union has size bigger than (${n\choose k}-{n-s\choose k}$.
For the proof of Theorem 1 we rely on the techniques from the paper of Kiselev and Kupavskii [12]. There, the authors studied colorings of $KG_{n,k}$ into $n-2k+2$ colors without trivial colors, and managed to show that such colorings do not exist for $n>(2+\epsilon)k^{2}$. The question of Erdős is related to the famous Erdős Matching Conjecture. In terms of $KG_{n,k}$, it asks for the largest subset of vertices with no $(s+1)$-clique. The extremal examples for $n>Csk$ are the same as for the problem discussed above, albeit the regimes are different: the EMC becomes trivial for $n<(s+1)k$. See the papers of Frankl, Kolupaev and the second author [6, 13] for the most recent results. It is also worth mentioning a recent paper of Frankl and the second author [7], in which they studied the size of the largest intersecting family in (${[n]\choose k}$ with large $\tau$ and showed that there is a certain ‘phase transition’ around $n=k^{2}$.
Another question that we address concerns the choice number of $KG_{n,k}.$ Recall that the choice number $ch(G)$ of a graph $G$ is the smallest $\ell$ such that, whenever each vertex of the graph is assigned a list of colors of size at least $\ell,$ we can choose one color from the list for each vertex so that the resulting coloring is proper (i.e., does not have monochromatic edges). Bulankina and Kupavskii [1] showed that $ch(KG_{n,k})\geq\frac{1}{2\ell^{2}}n\log n$ for $k\leq n^{\frac{1}{2}-\frac{1}{\ell}}$ and sufficiently large $n$. They also showed the upper bound $ch(KG_{n,k})\leq n\ln\frac{n}{k}+n,$ valid for any $n\geq 2k.$ In this paper, we improve their lower bound.
Theorem 3.
Fix $\varepsilon>0$ and let $n\geq n_{0}(\varepsilon)$. Suppose that $3\leq k\leq\left(\frac{n}{1+\varepsilon}\right)^{1/2}$. Then
$ch(KG_{n,k})>\left(\frac{1}{4}-\frac{1}{2k}-\varepsilon\right)\frac{n\log{n}}{2}.$
The proofs of Theorems 1 and 3 both rely on structural properties of intersecting families that were studied in the paper of Kiselev and the second author [12].
2 Proofs
2.1 Preliminaries
Using the Hilton–Milner theorem, it is easy to give a bound for Theorem 1 in the following regime.
Proposition 1.
Let $n>k^{2}+k^{3/2}\sqrt{s}$, $s\leq n$, and ($\mathcal{F}\subset{[n]\choose k}$ is a union of $s$ non-trivial intersecting families, then ||<($|\mathcal{F}|<{n\choose k}-{n-s\choose k}.$
We remark that it covers the case $n\geq 2k^{3}$.
Proof.
The Hilton-Milner Theorem implies
$|\mathcal{F}|\leq s\Big{(}{n-1\choose k-1}-{n-k-1\choose k-1}+1\Big{)}\leq sk{n-2\choose k-2}\leq\frac{sk^{3}}{n^{2}}{n\choose k}.$ (2.1)
We also have
(${n\choose k}-{n-s\choose k}={n\choose k}\Big{(}1-\prod_{i=0}^{k-1}\frac{n-s-i}{n-i}\Big{)}\geq{n\choose k}\Big{(}1-e^{-\sum_{i=0}^{k-1}\frac{s}{n-i}}\Big{)}\geq{n\choose k}\Big{(}1-e^{-\frac{sk}{n}}\Big{)}.$ (2.2)
The RHS of (2.1) is smaller than the RHS of (2.2) iff $1-\frac{sk^{3}}{n^{2}}>e^{-sk/n}.$ We have
$1-\frac{sk^{3}}{n^{2}}=\Big{(}1+\frac{sk^{3}}{n^{2}-sk^{3}}\Big{)}^{-1}\geq e^{\frac{-sk^{3}}{n^{2}-sk^{3}}}.$ (2.3)
Thus, the conclusion of the proposition holds, as long as
$\frac{sk^{3}}{n^{2}-sk^{3}}<\frac{sk}{n}$
We solve the quadratic equation $n^{2}-nk^{2}-sk^{3}>0$ and get $n>\frac{k^{2}}{2}(1+\sqrt{1+\frac{4s}{k}})$. For $n>k^{2}+k^{3/2}\sqrt{s}$ this inequality holds. ∎
We call families $\mathcal{A},\mathcal{B}\subset 2^{X}$ cross-intersecting, if $A\cap B\neq\emptyset$ for any $A\in\mathcal{A},B\in\mathcal{B}$. We say that $C\subset X$ covers family $\mathcal{F}\subset 2^{X}$ if $C\cap A\neq\emptyset$ for any $A\in\mathcal{F}$. We denote the size of the smallest cover of $\mathcal{F}$ by $\tau(\mathcal{F})$. We say that a family $\mathcal{H}$ set-covers a family $\mathcal{F}$ if for each $F\in\mathcal{F}$ there exists $H\in\mathcal{H}$ such that $H\subset F$. For a graph $G$ denote $\mathcal{I}_{k}(G)$ the family of independent $k$-element sets of vertices of $G$. In the proofs of Theorems 1 and 3, we need the following result from [12]:
Theorem 4 (Lemma 12 in [12]).
Let ($\mathcal{F}\subset{[n]\choose k}$ be an intersecting family with $\tau(\mathcal{F})\geq 2$. Then we can split it into $\mathcal{F}^{{}^{\prime}}\sqcup\mathcal{F}^{{}^{\prime\prime}}$, where $\mathcal{F}^{{}^{\prime}}$ can be set-covered by at most $k$ $2$-edges and $\mathcal{F}^{{}^{\prime\prime}}$ cross-intersects some family of $t$-sets $\mathcal{G}^{\times}$ with $t\in\{k-1,k\}$ and $\tau(\mathcal{G}^{\times})\geq\sqrt{k}$.
We can deduce the following structural property for the family $\mathcal{F}^{\prime\prime}$ from the theorem above.
Lemma 1.
Let $\mathcal{F}$ be a family of $k$-sets, which cross-intersects a family of $t$-sets $\mathcal{G}^{\times}$ with $t\in\{k-1,k\}$ and $\tau({\mathcal{G}^{\times}})\geq\sqrt{k}$. Then $\mathcal{F}$ can be set-covered by a family $\mathcal{H}$ of at most $k^{\sqrt{k}}$ $\sqrt{k}$-sets.
We construct $\mathcal{H}$ inductively. Put $\mathcal{H}_{0}=\emptyset$. For each $l\in[0,\sqrt{k}-1]$, we construct a set-cover $\mathcal{H}_{l+1}\subset{[n]\choose l+1}$ for $\mathcal{F}$ from the set-cover $\mathcal{H}_{l}\subset{[n]\choose l}$ that we assume to have constructed. We also assume that $|\mathcal{H}_{l}|\leq k^{l}$. Recall that $\mathcal{H}_{l}$ being a set-cover means that $\mathcal{F}\subset\bigcup_{H\in\mathcal{H}_{l}}\mathcal{F}(H)$. Thus, it should be clear that $\mathcal{H}_{0}$ satisfies both the condition on size and on being a set-cover. For each $H\in\mathcal{H}_{l}$ consider the set $G_{H}\in\mathcal{G}^{\times}$ disjoint with $H$. Such set exists since $|H|=l<\sqrt{k}$ and $\tau(\mathcal{G}^{\times})\geq\sqrt{k}$. Put $\mathcal{H}_{l+1}=\{H\cup x:H\in\mathcal{H}_{l},x\in G_{H}\}$. Then it is easy to see that $|\mathcal{H}_{l+1}|\leq t|\mathcal{H}_{l}|\leq k^{l+1}$ and
$\mathcal{F}\subset\bigcup_{H\in\mathcal{H}_{l+1}}\mathcal{F}(G).$
Finally, we put $\mathcal{H}=\mathcal{H}_{\sqrt{k}}$. ∎
Combining two previous statements, we get the following corollary.
Corollary 2.
Let $\mathcal{F}$ be an intersecting family of $k$-sets. If $\mathcal{F}$ is non-trivial then it is a subfamily of $\mathcal{A}\cup\mathcal{B}$, where the family $\mathcal{A}$ is the family of all $k$-sets that are set-covered by a family ($\mathcal{E}\subset{[n]\choose 2}$ with $|\mathcal{E}|=k,$ and the family $\mathcal{B}$ is the family of $k$-sets that are set-covered by a family ($\mathcal{H}\subset{[n]\choose\sqrt{k}}$ with $|\mathcal{H}|=k^{\sqrt{k}}.$ If $\mathcal{F}$ is trivial it is a subfamily of some star $\mathcal{S}_{x}$.
2.2 Proof of Theorem 1
In the proof of the theorem, we shall need the following result from [12].
Theorem 5 (Theorem 13 in [12]).
Let $n,k$ be sufficiently large integers such that $n\geq\big{(}2+\frac{5}{\ln k}\big{)}k^{2}$ and let $s<n$. Let $G$ be a graph on $[n]$ with at most $sk$ edges. Let $\mathcal{H}\subset\mathcal{I}_{k}(G)$ be a family of independent $k$-sets in $G$, which cross-intersects some family $\mathcal{G}$ of $t$-sets with $\tau(\mathcal{G})\geq\sqrt{k}$ and $t\leq k$. Then $|\mathcal{I}_{k}(G)|>e^{k^{0.1}/2}|\mathcal{H}|$.
We also need the following result from by Khadžiivanov and Nikiforov [9]:
Theorem 6.
[Khadžiivanov–Nikiforov] For a given graph $G$ let $\gamma$ be the density $\frac{|E(G)|}{|V(G)|^{2}}$ and $\mathcal{N}_{k}(G)$ be the number of cliques on $k$ vertices. Then, if $\gamma\geq\frac{k-2}{2(k-1)}$, we have
$\mathcal{N}_{k}(G)\geq\frac{1+(2\gamma-1)(k-1)}{k}\cdot|V(G)|\cdot\mathcal{N}_{k-1}(G)\text{ and }\mathcal{N}_{k-1}(G)>0.$
We prove Theorem 1 in assumption that $n<2k^{3}$. The case $n\geq 2k^{3}$ is covered by Proposition 1.
  • Proof of Theorem 1. We assume that $n<2k^{3}$ and thus that $k$ itself is sufficiently large. Let $\mathcal{F}_{1},\dots,\mathcal{F}_{s}$ be a collection of $s$ non-trivial intersecting families. By Theorem 4, we can split each $\mathcal{F}_{i}$ into $\mathcal{F}^{{}^{\prime}}_{i}\sqcup\mathcal{F}^{{}^{\prime\prime}}_{i}$, where $\mathcal{F}^{{}^{\prime}}_{i}$ can be set-covered by at most $k$ $2$-edges and $\mathcal{F}^{{}^{\prime\prime}}_{i}$ cross-intersects some family of $t$-sets $\mathcal{G}$ with $t\in\{k-1,k\}$ and $\tau(\mathcal{G})\geq\sqrt{k}$. We get a split $\cup_{i\in[s]}\mathcal{F}_{i}=\cup_{i\in[s]}(\mathcal{F}^{{}^{\prime}}_{i}\sqcup\mathcal{F}^{{}^{\prime\prime}}_{i})$. Next, we construct graph $G=([n],E),|E|\leq ks$, such that its set of edges $\mathcal{E}$ set-covers $\cup_{i\in[s]}\mathcal{F}^{{}^{\prime}}_{i}$. We denote by $\mathcal{I}_{k}(G)$ the family of independent $k$-sets of $G$. We may of course assume that $\cup_{i\in[s]}\mathcal{F}^{{}^{\prime\prime}}_{i}\subset\mathcal{I}_{k}(G)$.
    By Theorem 5, for any $\epsilon>0$, sufficiently large $k$ and $n>(2+\epsilon)k^{2}$ we have that $|\mathcal{I}_{k}(G)|>k^{4}|\mathcal{F}^{{}^{\prime\prime}}_{i}|$ for any $i\in[s]$. Hence,
    $\frac{s}{k^{4}}|\mathcal{I}_{k}(G)|>|\cup_{i\in[s]}\mathcal{F}^{{}^{\prime\prime}}_{i}|.$ (2.4)
    We need the following theorem, which we prove below.
    Theorem 7.
    Let $n,k$ be sufficiently large integers such that $2k^{2}<n<2k^{3}$. Let $s<n$. Let $G$ be a graph on $[n]$ with at most $sk$ edges and let $\mathcal{I}_{k}(G)$ be the family of all $k$-sets in $G$ that are independent in $G$. Then $|\mathcal{I}_{k}(G)|>\frac{k^{4}}{k^{4}-s}{n-s\choose k}$.
Using Theorem 7 and inequality (2.4), it is straightforward to finish the proof of Theorem 1.
$\Big{|}\bigcup_{i\in[s]}\mathcal{F}_{i}\Big{|}={n\choose k}-|\mathcal{I}_{k}(G)|+\Big{|}\bigcup_{i\in[s]}\mathcal{F}^{{}^{\prime\prime}}_{i}\Big{|}<{n\choose k}-\frac{k^{4}-s}{k^{4}-s}{n-s\choose k}={n\choose k}-{n-s\choose k}.$
Next, we prove Theorem 7.
  • Proof of Theorem 7. The quantity $|\mathcal{I}_{k}(G)|$ is equal to $\mathcal{N}_{k}(\bar{G})$, where $\bar{G}$ is the complement of $G$. We iteratively apply Theorem 6 in order to bound $\mathcal{N}_{k}(\bar{G})$:
    $\mathcal{N}_{k}(\bar{G})\geq\frac{2(k-1)\gamma-(k-2)}{k}|V(\bar{G})|\mathcal{N}_{k-1}(\bar{G}),$
    $\dots$
    $\mathcal{N}_{3}(\bar{G})\geq\frac{2\cdot 2\gamma-1}{3}|V(\bar{G})|\mathcal{N}_{2}(\bar{G}).$
    In order to apply Theorem 6, we need the inequality $\gamma\geq\frac{k-2}{2(k-1)}$ to hold. We have $\gamma=({n\choose 2}-ks)/n^{2}$. Rearranging, we get a quadratic inequality $n^{2}-(k-1)n-2k(k-1)s\geq 0$. Since $s<n$, we have $n^{2}-(k-1)n-2k(k-1)s>n(n-2k(k-1))-(k-1)n=n(n-(2k+1)(k-1))>0$ for $n>2k^{2}$. In what follows, we use $\mathcal{N}_{2}(\bar{G})=|E(\bar{G})|={{n\choose 2}-ks}$ and $2\gamma-1=-\frac{1}{n}-\frac{2ks}{n^{2}}$.
    $\mathcal{N}_{k}(\bar{G})\geq$ $n^{k-2}\left({n\choose 2}-ks\right)\prod_{i=2}^{k-1}\frac{1+(2\gamma-1)i}{i+1}$
    $=$ $\frac{n^{k}}{k!}\cdot\prod_{i=1}^{k-1}\left(1-i\left(\frac{1}{n}+\frac{2ks}{n^{2}}\right)\right).$
    In what follows, we assume that (${n-s\choose k}>0$ (equivalent to $n\geq s+k$), otherwise the statement is trivial. In order to prove the theorem, it is thus sufficient for us to show that
    $\prod_{i=1}^{k-1}\left(1-i\left(\frac{1}{n}+\frac{2ks}{n^{2}}\right)\right)\cdot\left(1-\frac{s}{k^{4}}\right)>\frac{k!{n-s\choose k}}{n^{k}}=\prod_{i=0}^{k-1}\left(1-\frac{s+i}{n}\right).$
    Since $n<2k^{3}\leq k^{4}$, we have $\big{(}1-\frac{s}{k^{4}}\big{)}>\big{(}1-\frac{s}{n}\big{)}$, and thus the above inequality is implied by
    $\prod_{i=1}^{k-1}\left(1-i\left(\frac{1}{n}+\frac{2ks}{n^{2}}\right)\right)>\prod_{i=1}^{k-1}\left(1-\frac{s+i}{n}\right).$
    We will simply show this term by term: for each $i\in[k-1]$, we have
    $1-i\left(\frac{1}{n}+\frac{2ks}{n^{2}}\right)>1-\frac{s+i}{n}.$
    This is equivalent to $sn>2kis$, or $n>2ki$, which is valid since $n>2k^{2}\geq 2ki$. ∎
2.3 Proof of Theorem 2
In order to prove the theorem, we give an explicit construction of a collection of non-trivial intersecting families whose union is bigger than that of the same number of trivial intersecting families.
Let $n,k$ be positive integers, $m$ a non-negative integer, $m\leq k-2$ and $n=(k-1)(k+1-m)+m$. Take the following pairwise disjoint sets: sets $T_{1},\ldots,T_{k-1}$ of size $3$; sets $B_{1},\ldots,B_{k-1}$ of size $k-2-m$; a set $M$ of size $m$. The ground set for the construction is the union of all these sets:
$G=M\sqcup\bigsqcup_{i=1}^{k-1}(T_{i}\sqcup B_{i}).$
Note that $|G|=n$. The majority of the families in the construction would be of Hilton-Milner type. In order to define them, we need to define the ‘exceptional’ sets. For each $i\in[k-1]$ and $y\in B_{i}$ put
$H^{i}_{y}:=M\cup T_{i}\cup B_{i}\setminus\{y\}.$
Consider the following collection of intersecting families (cf. Figure 1). For each $i\in[k-1]$ and $y\in B_{i}$, define
$\mathcal{HM}^{i}_{y}=\{H_{y}^{i}\}\cup\Big{\{}A\in{G\choose k}:y\in A,A\cap H_{y}^{i}\neq\emptyset\Big{\}}.$
Note that there are $\sum_{i=1}^{k-1}|B_{i}|=(k-1)(k-2-m)$ such families. Next, for each $i=1,\ldots,k-1$ define a family $\mathcal{T}_{i}$ of the form
$\mathcal{T}_{i}:=\Big{\{}A\in{G\choose k}:|A\cap T_{i}|\geq 2\Big{\}}.$
[…]
Figure 1: Construction of intersecting families whose union is bigger than the union of stars. Red boxes correspond to the parts of the ground set. Their sizes are indicated on the right. Green ellipses are the 2-element sets that define $\mathcal{T}_{i}$. Blue vertex $y$ and a set $H_{y}^{i}$ define $\mathcal{HM}_{y}^{i}$.
It is not difficult to see that the families $\mathcal{HM}_{y}^{i}$ and $\mathcal{T}_{i}$ are all non-trivial intersecting, moreover, their total number is $(k-1)(k-2-m)+k-1=n-2k+2-m.$
Let us determine, how many sets $A\in{G\choose k}$ are not covered by these families. If $|A\cap T_{i}|\geq 2$ then $A$ is covered by $\mathcal{T}_{i}$. $y\in A\cap B_{i}$ and $|A\cap(T_{i}\sqcup B_{i})|\geq 2$ then $A\in\mathcal{HM}_{y}^{i}$. Thus, all $A$ with $|A\cap(T_{i}\sqcup B_{i})|\geq 2$ are covered by our families.
If $A$ is such that $|A\cap(T_{i}\sqcup B_{i})|\leq 1$ for all $i\in[k-1]$, then, by the pigeon-hole principle, $|A\cap M|\geq 1$. If at the same time $y\in A\cap B_{i}$ for some $i$ then $A\in\mathcal{HM}_{y}^{i}$.
Concluding, $A$ is not covered by these families if (and only if) $A\subset M\sqcup\bigsqcup_{i=1}^{k-1}T_{i}$ and $|A\cap T_{i}|\leq 1$ for each $i\in[k-1]$. The number of such sets is
$\sum_{j=1}^{m}{m\choose j}{k-1\choose k-j}3^{k-j}.$
For $m=1$ this sum is equal to $3^{k-1}$. On the other hand, the number of sets not covered by a union of $n-2k+2-m$ stars is
(${2k-2+m\choose k}.$
For $m=1$, it is equal to $(4+o(1))^{k}$. It is not difficult, although tedious, to check that the first expression is smaller than the second for large enough $k$ and $m\leq k/4$. We decided to omit these calculations.
2.4 Proof of Theorem 3
We employ the probabilistic method. Put $u:=Cn\log{n}$ with some $C$ that we specify later on. Take a set of $u$ colors and correspond to each vertex of $KG_{n,k}$ a random subset of colors of size $\frac{u}{2}$. Any independent set of vertices of $KG_{n,k}$ is an intersecting family in (${[n]\choose k}$. Using Corollary 2, we get that any such independent set is contained in one of the families from $\mathcal{C}$, where $\mathcal{C}$ consists of families of the form $\mathcal{A}\cup\mathcal{B}$ or $\mathcal{S}_{x}$. Note that
$|\mathcal{C}|\leq{{n\choose 2}\choose k}{{n\choose\sqrt{k}}\choose k^{\sqrt{k}}}+n.$
We say that a partition $X=X_{1}\sqcup\dots\sqcup X_{m}$ of a set $X$ lies in a cover of this set $X=X^{\prime}_{1}\cup\dots\cup X^{\prime}_{m}$, if $X_{i}\subset X^{\prime}_{i}$ for each $i\in[m]$. Hence, any partition of (${[n]\choose k}$ lies in one of
$|\mathcal{C}|^{u}\leq\left({{n\choose 2}\choose k}{{n\choose\sqrt{k}}\choose k^{\sqrt{k}}}+n\right)^{u}\leq n^{\left(2k+k^{\sqrt{k}+\frac{1}{2}}\right)u}$
possible covers, formed by a $u$-tuple of families from $\mathcal{C}$.
For a fixed cover $\mathcal{K}=\mathcal{K}_{1}\cup\dots\cup\mathcal{K}_{u}$ from $\mathcal{C}$ we are going to bound from above the probability of the event $\mathcal{A_{K}}$ that $KG_{n,k}$ can be colored in one of the colorings of (${[n]\choose k}$ that lies in $\mathcal{K}$ and that respects the lists of colors assigned to the $k$-sets.
First, we show that families of type $\mathcal{B}$ cover only a small part of vertices of $KG_{n,k}$. Consider an arbitrary family of type $\mathcal{B}$. Clearly, it covers less than $k^{\sqrt{k}}{n-\sqrt{k}\choose k-\sqrt{k}}$ $k$-sets. We have at most $u$ families of such type, and thus together they can cover at most
$uk^{\sqrt{k}}{n-\sqrt{k}\choose k-\sqrt{k}}\leq Cn\log{n}\>k^{\sqrt{k}}\left(\frac{k}{n}\right)^{\sqrt{k}}{n\choose k}$
sets. We have
$\frac{Cn\log{n}\>k^{2\sqrt{k}}}{n^{\sqrt{k}}}<\frac{1}{6}$
for large enough $n$ that also satisfies $n>(1+\varepsilon)k^{2}$. Therefore, the number of $k$-sets covered by families of type $\mathcal{B}$ is less than $\frac{1}{6}{n\choose k}$.
Second, we analyze, what could be covered by the families $\mathcal{A}$ and $\mathcal{S}_{x}$. Consider a random $k$-set $S$ from $[n]$. Let $\eta_{1}$ denote the number of $x\in S$ such that $x$ is a center of some $\mathcal{S}_{x}$. Let $\eta_{2}$ denote the number of pairs $\{x,y\}\subset S$ such that $\{x,y\}$ is contained in $\mathcal{E}(\mathcal{A})$ for some $\mathcal{A}.$ For convenience, we also call such pairs centers. Then the random variable $\eta_{1}+\eta_{2}$ equals to the number of centers of type $\mathcal{S}_{x}$ or $\mathcal{A}$ in a random $k$-set.
Suppose that there are $u_{1}$ families of type $\mathcal{S}_{x}$ and $u_{2}$ families of type $\mathcal{A}$ in the cover $\mathcal{K}$. Using that each family of type $\mathcal{A}$ consists of at most $k$ pairs and that $n\geq k^{2}$, we can upper bound the expectation of $\eta_{1}+\eta_{2}$:
$\operatorname{\mathrm{E}}[\eta_{1}+\eta_{2}]\leq u_{1}\frac{k}{n}+u_{2}\frac{k{k\choose 2}}{{n\choose 2}}\leq u_{1}\frac{k}{n}+u_{2}\frac{k^{3}}{n^{2}}\leq u_{1}k/n+u_{2}k/n\leq uk/n.$
Using Markov’s inequality, we get that
Pr[ a uniformly random set from ([$\Pr\left[\text{ a uniformly random set from }{[n]\choose k}\text{ contains more than }\frac{2uk}{n}\text{ centers }\right]\leq\frac{1}{2}.$
Thus, at least half of all (${[n]\choose k}$ sets contain at most $\frac{2uk}{n}$ centers. It means that each of such sets can be colored in at most $\frac{2uk}{n}$ colors of type $\mathcal{S}_{x}$ or $\mathcal{A}$. (At this point, we ignore the color lists assigned to the vertices.) From the previous paragraph, we also know that most of them cannot be colored in colors of type $\mathcal{B}$. Let $\mathcal{X}$ denote the family of sets which can only be colored in at most $\frac{2uk}{n}$ colors of types $\mathcal{S}_{x}$ or $\mathcal{A}$. Then
$|\mathcal{X}|\geq\frac{1}{3}{n\choose k}.$
Given $X\in\mathcal{X}$ and the color list assigned to $X$, let $B_{X,\mathcal{K}}$ stand for the event that $X$ cannot be colored with a coloring that lies in $\mathcal{K}$. $X$ can be colored in color $j$, where $\mathcal{K}_{j}$ of type $\mathcal{S}_{x}$ or $\mathcal{A}$ only if $j$ belongs to the randomly chosen subset of $\frac{u}{2}$ colors that was assigned to $\mathcal{X}$. There are $2uk/n=2Ck\log_{2}n=:z$ colors that contain $X$. Then the probability $p$ of the event $B_{X,\mathcal{K}}$ is at least the probability that the colors for $X$ were chosen from the complement of the set of possible colors:
$p\geq\frac{{u-z\choose u/2}}{{u\choose u/2}}=2^{-z}(1-o(1))\geq\frac{1}{2}n^{-2Ck}.$
Recall that $\mathcal{A_{K}}$ is the event that $KG_{n,k}$ can be colored in one of the colorings of (${[n]\choose k}$ that lies in $\mathcal{K}$ and that respects the lists of colors assigned to the $k$-sets. We have $\mathcal{A}_{\mathcal{K}}\subset\bigcap_{X\in\mathcal{X}}\bar{B}_{X,\mathcal{K}}$, and thus
$\Pr[\mathcal{A}_{\mathcal{K}}]\leq(1-p)^{|\mathcal{X}|}\leq e^{-p|\mathcal{X}|}.$
We have
$p|\mathcal{X}|\geq\frac{1}{6}n^{-2Ck}{n\choose k}\geq\frac{1}{6}n^{-2Ck}\left(\frac{n}{k}\right)^{k}\geq n^{k/2-2Ck},$
for large enough $n>(1+\varepsilon)k^{2}$, therefore
$\Pr[\mathcal{A}_{\mathcal{K}}]\leq e^{-n^{k/2-2Ck}}.$
Let $U$ denote the event that there exists a proper coloring of $KG_{n,k}$ that respects the lists assigned to the vertices.
To conclude the proof, it is enough to verify the last inequality below.
$\Pr[U]\leq\sum_{\mathcal{K}\in\mathcal{C}^{u}}\Pr[\mathcal{A}_{\mathcal{K}}]\leq n^{\left(2k+k^{\sqrt{k}+\frac{1}{2}}\right)u}\cdot e^{-n^{k/2-2Ck}}<1.$
Comparing the exponents, it is equivalent to
$2k+k^{\sqrt{k}+1/2}<\frac{n^{k/2-1-2Ck}}{C\log^{2}{n}}.$
The RHS is at least $2n^{k(1/2-2C)-1-\delta}$ for an arbitrary $\delta>0$ and $n\geq n_{0}(\delta)$, and the LHS is at most $2k^{\sqrt{k}+1/2}$. Thus, it is sufficient to check the inequality
$k^{\sqrt{k}+1/2}<n^{k(1/2-2C)-1-\delta}.$ (2.5)
It is easy to check that it holds if $1/2-2C\geq\delta+\frac{1}{k}$ and $n$ sufficiently large. The inequality we get on $C$ is $C\leq\frac{1}{4}-\frac{1}{2k}-\epsilon.$ For $k=3$, for example, this gives $\frac{1}{12}-\epsilon.$ The theorem is proved.